An 80-kg firefighter slides down a brass pole 8.0 m high in a fire station while exerting a frictional force of 600 N on the pole in the time of?



1
Expert's answer
2021-10-24T18:28:08-0400

"ma=mg-F_{fr}"

"mx''(t)=mg-F_{fr}"

"x=(g-F_{fr}\/m)t^2\/2+c_1t+c_2"

if "v_0=0" then "c_1=0,c_2=0"


So:

"t=\sqrt{\frac{2\cdot8}{g-F_{fr}/m}}=\sqrt{\frac{16}{9.8-600/80}}=2.6" s